OpenStudy (anonymous):

Divide the following polynomials, then place the answer in the proper location on the grid. Write the answer in descending powers of x. 9x 2 - 18x - 7 ÷ (3x + 1)

OpenStudy (campbell_st):

well the quadratic can be factored it makes the division easier for a quadratic \[ax^2 + bx +c\] multiply a and c so in your question a = 9 and c = -7 so a*c = -63 find the factors of -63 that add to -18 the coefficient of b the larger factor will be negative then use \[\frac{(ax + factor ~1)(ax + factor ~2)}{a}\] or in your question \[\frac{(9x + factor~1)(9x + factor ~2)}{9}\] then remove common factors in the numerator... and you'll find they cancel with the denominator